OpenStudy (anonymous):

What is the approximate volume of the cone? Use 3.14 for π. A. 6359 cm3 B. 1272 cm3 C. 2120 cm3 D. 4239 cm3

OpenStudy (anonymous):

@LiteLegacy @rvc

OpenStudy (anonymous):

@khalilforthewin

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

\(\sf V = \dfrac{1}{3} \pi r^2\)

OpenStudy (igreen):

Oops, left the height out.

OpenStudy (igreen):

It should be: \(\sf V = \dfrac{1}{3} \pi r^2 h\) Anyway, Plug in what we know: \(\sf V = \dfrac{1}{3}(3.14)(15^2)(9)\)

OpenStudy (igreen):

Can you simplify that?

OpenStudy (anonymous):

ill try

OpenStudy (igreen):

Okay, simplify the exponent first.

OpenStudy (igreen):

\(\sf 15^2 \rightarrow 15 \times 15\)

OpenStudy (anonymous):

225

OpenStudy (igreen):

Yes, now multiply it with the rest. \(\sf V = \dfrac{1}{3}(3.14)(225)(9)\)

OpenStudy (anonymous):

2119.5

OpenStudy (igreen):

Yep, that's our answer.

OpenStudy (anonymous):

thank you

OpenStudy (unklerhaukus):

don't forget units

OpenStudy (igreen):

Yes, when dealing with 3-D objects, our regular centimeters becomes cubic centimeters(\(\sf cm^3\)).